# Ada-NAV: Adaptive Trajectory Length-Based Sample Efficient Policy Learning for Robotic Navigation

### v4

@@ III-A Key Insight: Policy Entropy as a Proxy for Mixing Time
Given a policy $\pi_{\theta}$, we would ideally like to base trajectory length for our PG updates on the spectral gap from the mixing time lower bound in Theorem 4. However, without prior knowledge of the transition kernel $\mathbb{P}$ of the underlying MDP, it is difficult to calculate the spectral gap of the induced transition matrix, $\mathbb{P}_{\pi_{\theta}}$. One possible workaround is to estimate the transition kernel, but this quickly becomes intractable due to the $\Omega(|\mathcal{S}||\mathcal{A}|)$ number of samples required for each estimate. Interestingly, one key insight we observe is that if the policy becomes more deterministic and thus the entropy decreases, the spectral gap of the induced transition matrix also decreases or stays the same, suggesting a monotonic relationship between mixing time and policy determinism.
Figure 2 illustrates this correlation with a 2D gridworld environment. We have a $25$-by-$25$ gridworld. Each plot in Figure 2 corresponds to a different fixed transition probability kernel $\mathbb{P}$. For each fixed kernel, we construct 100 tabular policies, each with a discrete action space of 5 elements to represent up, down, left, right, and stay. The tabular policies have various average entropies, ranging from maximum entropy with uniform probability over the action space for each state (end of x-axis on Figure 2) to completely deterministic for each state (beginning of x-axis) with a probability of $1$ of moving upwards. We define policy entropy in for tabular policies as the average Shannon entropy across all states, see Equation 5. For each policy, we multiply the policy with the fixed transition kernel to calculate the underlying Markov Chain and then estimate its spectral gap.
In this environment, if the agent chooses an action that would make it go off the edge, the agent will stay in the current state. Both transition kernels are deterministic: each state-action pair is mapped to a single next state a potential agent can occupy. One kernel represents an empty gridworld with no obstacles. The relationship between spectral gap and policy entropy is plotted in orange for this kernel. The other kernel is represents a gridworld with $4$ walls dividing it into $4$ equal regions. Please see Appendix for visualization. The plot corresponding to this kernel is blue in Figure 2.  [^1] Code to notebook to generate plot: https://github.com/Bridge00/spectral_gap_policy_entropy.git
To be clear, in these experiments, we have not trained any agent to navigate in the gridworld. We have a static environment where we manually designed transition kernel matrices and multiplied each one with 100 different tabular policies to estimate the underlying Markov Chains and their corresponding spectral gaps, a task normally intractable for more complex, realistic environments. For results displaying Ada-NAV’s effectiveness in increasing sample efficiency for a goal-reaching task in this environment and other gridworld environments, please see the Appendix.
Figure 2: Positive correlation between a policy’s entropy and the spectral gap of its induced transition matrix over a $25$-by-$25$ gridworld environment with different fixed transition kernels. Policy entropy is a viable proxy for mixing time.
@@
where $N$ is the action space dimension and $\Sigma_{\theta}(s)$ is the covariance matrix of the policy $\pi_{\theta}(\cdot|s)$. We note that there are no analytical equations to calculate mixing time, and calculating spectral gap requires knowledge of the transition dynamics $P(s^{\prime}|s,a)$ which restricts it to only model-based RL. Thus, using analytical Equations (5) and (6) as a proxy for mixing time is a more simple method that can be used for either model-free or model-based RL.
III-B Ada-NAV: Mapping Policy Entropy to Trajectory Length
@@
We now present our novel adaptive trajectory length scheme for robotic navigation. The key to our approach is to select trajectory length based on the current policy entropy value. We perform this computation using a monotonic mapping from the range of possible policy entropy values to a range of possible trajectory length values. Let $H_{i}$ denote the maximum possible policy entropy value and $H_{c}$ the policy entropy of the current policy. In practice, we choose $H_{i}=H(\pi_{\theta_{0}})$ to be the entropy of the initial policy, since policy entropy tends to decrease during training. Let $t_{i}\in\mathbb{N}$ denote the user-specified initial trajectory lengths and let $\eta>0$ denote a user-specified sensitivity parameter that controls the rate of change of length (see the Appendix for specific values of $t_{i},\eta$ used). For our experiments, we determine the trajectory length corresponding to the current policy entropy using the mapping $f$ below:
$t_{c}=f^{H_{i}}_{t_{i},\eta}(H_{c})=\text{round}\left(1-\frac{H_{c}}{H_{i}}(1-\frac{t_{i}}{\eta})\right),$ (7)
@@
where the rounding operation is to ensure that $t_{c}$ is always an integer. In the Appendix, we utilize other monotonic transformation to map entropy to trajectory length in the form of an exponential function. Furthermore, in our experiments, we initialize our policies with a Gaussian distribution. To ensure, that $H_{c}$ never increases too much so that $t_{c}$ becomes 00, one can initialize with a uniform distribution so that $H_{i}$ is the maximum possible entropy. Algorithm 1 summarizes our Ada-NAV methodology for adaptive trajectory length PG methods. Note that any on-policy PG method can used to carry out the policy gradient update in line 3. The training is divided into episodes where the start location and goal are constant between episodes. During each episode, a gradient update occurs after every trajectory. The amount of samples in each trajectory is the trajectory length. The episode ends when goal is reached or when a set number of interactions have occurred. For clarification, given an episode, the trajectory length is fixed. At the start of the new episode is when the trajectory length is adjusted based on policy entropy.
@@
1:  Initialize: policy parameter $\theta_{0}$, policy entropy $H_{i}$, trajectory length $t_{i}$, current trajectory length $t_{c}=t_{i}$, maximum trajectory length $t_{d}$, number of episodes $E$, maximum number $K$ of iterations per episode, sensitivity parameter $\eta$, $t=0$
@@
3:     calculate $H_{c}$ using 5 for discrete actions, 6 for continuous actions
4:     calculate $t_{c}$ using 7
5:     $k=0$
6:     while $k<K$ do
@@
8:        Perform gradient update using $\mathcal{T}$ to obtain $\theta_{t+1}$
9:        $k=k+t_{c}$
10:        $t=t+1$
@@
13:  Return: $\pi_{\theta_{t}}$
Algorithm 1 Policy Learning with Ada-NAV
@@ IV Experiments and Results
This section provides a detailed empirical study and comparison of the proposed adaptive trajectory methods in Algorithm 1.
@@
Figure 3: Comparison of four different trajectory-length schemes on [LEFT] even and [RIGHT] uneven terrain navigation tasks in Unity-based outdoor robot simulator. For REINFORCE, PPO, and SAC, using Ada-NAV as the trajectory length scheme leads to a navigation policy with a higher cumulative reward return than using constant trajectory lengths. The drawn trajectories are sample representations of the robot’s odometry data collected during each navigation task. The trajectories pictured are representative of all three algorithms, since we found that each trajectory-length scheme leads to similar navigation behavior regardless of which RL algorithm is used. See Figures 4 and 6 for training details.
@@
Figure 4: [TOP] Learning curves during training and [BOTTOM] sample complexity savings during navigation tasks for the even terrain shown on the left of Figure 3. Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths. For each algorithm, Ada-NAV converges to a higher cumulative reward while expending less of the total budget of $\sim 4.5e4$ samples. We ran $8$ to $10$ independent replications for each of the 12 experiments. Although fixed trajectory length 200 performs comparably to Ada-NAV in terms of sample complexity, it achieves lower reward in all cases.
@@
Figure 5: Sample navigation trajectories generated in simulated and real-world environments by policies trained using Ada-NAV vs. fixed trajectory lengths at a fixed sample budget of $\sim 4.5e4$. We observe that the fixed trajectory length policies often fail at test time due to inefficient use of the available sample budget during training. In contrast, Ada-NAV policies successfully complete the navigation tasks after training with the same sample budget. Further, the experiments with a real Clearpath Husky robot validate that our method can be transferred onto real robotic systems without significant performance degradation. The drawn trajectories are sample representations of the robot’s odometry data collected during each task.
@@
RL Algorithms. We choose three RL algorithms to integrate Ada-NAV into for improved sample efficiency: REINFORCE, PPO, and SAC. As REINFORCE and PPO are both on-policy policy gradient methods and SAC is an off-policy method, we show that Ada-NAV can improve the sample efficiency of various classes of policy gradient algorithms.
@@
Baselines. Since our proposed approach relies on adaptive trajectory length, for each algorithm we test Ada-NAV with, we compare it against the same algorithm with fixed trajectory lengths. We also compare with fixed trajectory lengths with max-entropy regularizer as that also encourages exploration.
@@
We also incorporate a classical navigation algorithm Ego-graph [48], to highlight the navigation performance of our method under challenging terrain conditions. Ego-graph is a baseline method in [48] that utilizes robot-centric elevation data to obtain the actions that minimize the elevation gradient cost.
Figure 6: [TOP] Learning curves during training and [BOTTOM] sample complexity savings during navigation tasks for the uneven terrain shown on the right of Figure 3. Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths. For each algorithm, Ada-NAV converges to a higher cumulative reward while expending less of the total budget of $\sim 4.5e4$ samples. We ran $8$ to $10$ independent replications for each of the 12 experiments. Although fixed trajectory length 200 performs comparably to Ada-NAV in terms of sample complexity, it achieves lower reward in all cases.
@@
Environments: We utilize two environments in a unity-based outdoor robot simulator. The policies for the outdoor robot simulator are based on the REINFORCE algorithm [27]. Finally, we deploy navigation policies trained for a fixed sample budget using the robot simulator on a real Clearpath Husky robot. We present the details of the training environments below:
@@
Even and Uneven Terrain Navigation in Robotic Simulations: We use a Unity-based outdoor simulator with a Clearpath Husky robot model to train policies for two navigation tasks: 1. Goal Reaching, 2. Uneven terrain navigation using the Reinforce algorithms. The Unity simulator includes diverse terrains and elevations for training and testing. We restrict ourselves to a $100m\times 100m$ region during training. We used identical policy distribution parameters and neural network architectures with all four algorithms during training and evaluation. The policies are implemented using Pytorch in to train in the simulator equipped with a Clearpath Husky robot model, and a ROS Melodic platform. The training and simulations are executed on a workstation with an Intel Xeon 3.6 GHz CPU and an Nvidia Titan GPU.
@@
Our agent is a differential drive robot with a two-dimensional continuous action space $a=(v,\omega)$ (i.e. linear and angular velocities). The state space varies for the two navigation tasks and is obtained in real time from the simulated robot’s odometry and IMU sensors. The state inputs used in the two scenarios are, The simulator includes even and uneven outdoor terrains and a differential drive robot model with a two-dimensional continuous action space $a=(v,\omega)$ (i.e. linear and angular velocities). We perform two navigation tasks: 1. Even terrain, 2. Uneven terrain navigation using the robot’s local sensor data as state observations. The state space varies for the two navigation tasks and is obtained in real-time from the simulated robot’s odometry and IMU sensors. The state inputs used in the two scenarios are,
@@
  • $d_{goal}\in[0,1]$ Normalized distance (w.r.t the initial straight-line distance) between the robot and its goal;
  • $\alpha_{goal}\in[0,\pi]$ Angle between the robot’s heading direction and the goal;
  • $(\phi,\theta)\in[0,\pi]$ Roll and pitch angle of the robot respectively.
@@
where, $\mathcal{N}$ denotes Gaussian distribution with variance $\sigma$ and $\beta$ is the amplitude parameter. We set the variance $\sigma=0.2$ to ensure that the $r_{dist}$ only includes two narrow reward peaks near the goal and half-way from the goal. Similarly, $r_{elev}$ is defined as $r_{elev}=\mathbb{I}_{\{|\phi|\geq\pi/6\}}\cup\mathbb{I}_{\{|\theta|\geq\pi/6\}}$. We observe that our reward definitions are significantly sparse compared to the traditional dense rewards used in the literature[31].
(a) Even terrain navigation: The robot is placed in an obstacle-free even terrain environment and random goals in the range of 10-15 meters away from the starting position are given at each episode. The state space is $s=[d_{goal},\alpha_{goal}]$ and the overall reward formulation $r_{even}=r_{goal}+r_{head}$.
(b) Uneven terrain navigation: The robot is placed in an obstacle-free, highly uneven terrain environment. The state space is $s=[d_{goal},\alpha_{goal},\phi,\theta]$. The objective is to navigate the robot to a goal location along relatively even terrain regions to avoid possible robot flip-overs. The overall reward formulation $r_{uneven}=r_{goal}+r_{head}+r_{elev}$.
(c) Real-world Experiments: The real and simulated testing were conducted in different outdoor scenarios. Particularly, the real-world environments are reasonably different in terms of terrain structure (i.e. elevation gradient) and surface properties (i.e grass and tiny gravel regions with different levels of friction) which are not included or modeled in the simulator (e.g. elevation gain in the simulator is up to $\sim 4m$, however, we restricted it to $\sim 1.5m$ elevation gain in the real world for robot’s safety). Our Clearpath Husky robot is equipped with a VLP16 LiDAR and a laptop with an Intel i9 CPU and an Nvidia RTX 2080 GPU.
@@
Ada-NAV(Ours) 72 16.28 1.985
Table I: Navigation performance of the policies trained for even and uneven terrain navigation tasks at a fixed sample budget. SR denotes success rate, PL denotes average path length, and EC denotes elevation cost. Each method is tested for 100 trials in the outdoor simulator and the average values are reported in the table. Test environments are different from training environments.

### v7

@@ III-A Key Insight: Policy Entropy as a Proxy for Mixing Time
Given a policy $\pi_{\theta}$, we would ideally like to base trajectory length for our PG updates on the spectral gap $1-\lambda_{\theta}^{2}$ that lower bounds $\tau_{mix}^{\theta}$. However, without prior knowledge of the transition kernel $\mathbb{P}$ of the underlying MDP, it is difficult to calculate the spectral gap of the induced transition matrix, $\mathbb{P}_{\pi_{\theta}}$. One possible workaround is to estimate the transition kernel, but this quickly becomes intractable due to the $\Omega(|\mathcal{S}||\mathcal{A}|)$ number of samples required for each estimate.
Interestingly, one key insight we observe in $25$-by-$25$ gridworld environments is that if the policy entropy decreases, the spectral gap of the induced transition matrix also decreases or stays the same (Figure 2).
Remark: We have not trained any agent to navigate in the gridworld. We have a static environment where we manually designed transition kernel matrices and multiplied each one with $100$ different tabular policies with various entropies, from uniform distributions to deterministic, distributions to estimate the underlying Markov Chains and their corresponding spectral gaps. Because of the impracticality of calculating spectral gaps from more complex environments, we provide these proof-of-concept experiments with gridworld environments to support the formulation of our CCE approach.
Figure 2: [LEFT] Visualization of $25$-by-$25$ gridworld with $4$ walls (black) dividing it into equal regions. [RIGHT] Positive correlation between a policy’s entropy and the spectral gap of its induced transition matrix over two different gridworld environments. Policy entropy is a viable proxy for mixing time.
@@
where $N$ is the action space dimension and $\Sigma_{\theta}(s)$ is the covariance matrix of the policy $\pi_{\theta}(\cdot|s)$. We note that there are no analytical equations to calculate mixing time, and calculating spectral gap requires knowledge of the transition dynamics $P(s^{\prime}|s,a)$ in the form of a matrix, which is infeasible for complex environments. Thus, using analytical Equations (4) and (5) as a proxy for mixing time is a simple method that can be used for model-free or model-based RL.
III-B CCE: Confidence-Controlled Exploration
@@
We now present our novel CCE approach for robotic navigation. The key to our approach is to select trajectory length based on the current value of the policy entropy, our measure for confidence. We perform this computation using a monotonic mapping $f$ from the range of possible policy entropy values to a range of possible trajectory length values. In this paper, we utilize a linear function; however, we note that $f$ can be implemented with any monotonic mapping such as an exponential function. Let $H_{i}$ denote the maximum possible policy entropy value and $H_{c}$ the policy entropy of the current policy. In practice, we choose $H_{i}=H(\pi_{\theta_{0}})$ to be the entropy of the initial policy, since policy entropy tends to decrease during training. Let $K_{i}\in\mathbb{N}$ denote the user-specified initial trajectory lengths and let $\alpha>0$ denote a user-specified sensitivity parameter that controls the rate of change of length. For our experiments, we determine the trajectory length, $K_{c}$, corresponding to $H_{c}$ using the mapping $f$ below:
$K_{c}=f^{H_{i}}_{K_{i},\alpha}(H_{c})=\text{round}\left(\alpha-\frac{H_{c}}{H_{i}}(\alpha-K_{i})\right),$ (6)
@@
where the rounding operation is to ensure that $K_{c}$ is always an integer. Furthermore, in our experiments, we initialize our policies with a Gaussian distribution. To ensure, that $H_{c}$ never increases too much so that $K_{c}$ becomes 00, one can initialize with a uniform distribution so that $H_{i}$ is the maximum possible entropy. Algorithm 1 summarizes our CCE methodology for adaptive trajectory length PG methods. Note that any on-policy PG method can used to carry out the policy gradient update in line 3. The training is divided into episodes where the start location and goal are constant between episodes. During each episode, a gradient update occurs after every trajectory. The episode ends when the goal is reached or when a user-set number $T$ of interactions has occurred. For clarification, given an episode, the trajectory length is fixed. $K_{c}$ is only adjusted based on the policy entropy at the start of the new episode.
@@
1:  Initialize: policy parameter $\theta_{0}$, policy entropy $H_{i}$, trajectory length $K_{i}$, current trajectory length $K_{c}=t_{i}$, number of episodes $E$, maximum number $T$ of iterations per episode, sensitivity parameter $\alpha$, update index $u=0$
@@
3:     calculate $H_{c}$ using 4 for discrete actions, 5 for continuous actions
4:     calculate $K_{c}$ using 6
5:     $t=0$
6:     while $t<T$ do
@@
8:        Perform gradient update using $\mathcal{T}$ to obtain $\theta_{u+1}$
9:        $t=t+K_{c}$
10:        $u=u+1$
@@
13:  Return: $\pi_{\theta_{u}}$
Algorithm 1 Policy Learning with CCE
@@ IV Experiments and Results
This section provides a detailed empirical study and comparison of the proposed CCE method in Algorithm 1.
@@
Figure 3: Comparison of four different trajectory-length schemes on [LEFT] even and [RIGHT] uneven terrain navigation tasks in Unity-based outdoor robot simulator. For REINFORCE, PPO, and SAC, using CCE as the trajectory length scheme leads to a navigation policy with a shorter path length return than baselines with limited samples. The drawn trajectories are sample representations of the robot’s odometry data collected. The trajectories pictured are representative of all $3$ algorithms since we found that each trajectory-length scheme leads to similar navigation regardless of the RL algorithm. See Figure 4 for training details.
@@
Figure 4: [TOP] Learning curves during training and [BOTTOM] percent of sample budget of $\sim 4.5e4$ samples during navigation tasks for the even terrain shown on the left of Figure 3. Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths. For each algorithm, CCE converges to a higher cumulative reward while expending less of the total budget. We ran $8$ to $10$ independent replications for each of the $12$ experiments. Although fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing the episode to end.
@@
Figure 5: Sample navigation trajectories generated in simulated and real-world environments by policies trained with REINFORCE using CCE vs. fixed trajectory lengths at a fixed sample budget of $\sim 4.5e4$. We observed that the fixed trajectory length policies often fail at test time due to inefficient use of the available sample budget during training. In contrast, CCE policies successfully complete the tasks after training with the same limited sample budget. The experiments with a real Clearpath Husky robot validate that CCE can be transferred onto real robotic systems without significant performance degradation. The drawn trajectories are sample representations of the robot’s odometry data collected during each task.
@@
RL Algorithms. We choose three RL algorithms to integrate CCE into for improved sample efficiency: REINFORCE, PPO, and SAC. As REINFORCE and PPO are both on-policy policy gradient methods and SAC is an off-policy method, we show that CCE can improve the sample efficiency of various classes of policy gradient algorithms. For PPO and SAC, we adaptively change the size of the replay buffer. For information on hyperparameters, such as network architecture, $K_{i}$, and $\alpha$, please see Table II in the Appendix 111we re-provide the link to the supplemental version for convenience https://arxiv.org/abs/2306.06192 of the supplemental version.
@@
Baselines. Since our proposed approach relies on adaptive trajectory length, for each algorithm we test CCE with, we compare it against the same algorithm with fixed trajectory lengths. We also compare with fixed trajectory lengths with max-entropy regularizer as that encourages exploration.
@@
We also incorporate a classical navigation algorithm Ego-graph [51], to highlight the navigation performance of our method under challenging terrain conditions. Ego-graph is a baseline method in [51] that utilizes elevation data to obtain the actions that minimize the elevation gradient cost.
Figure 6: Similar to Figure 4 but for uneven terrain shown on the right of Figure 3. CCE still consistently reaches the highest cumulative reward while saving the largest percentage of training data. As in Figure 4, although fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing the episode to end.
@@
Environments: We utilize two environments in a unity-based outdoor robot simulator. Finally, we deploy navigation policies trained for a fixed sample budget using the robot simulator on a real Clearpath Husky robot. We present the details of the training environments below:
@@
Simulated Even and Uneven Terrain Navigation: We use a Unity-based outdoor simulator with a Clearpath Husky robot model to train policies for two navigation tasks: 1. Goal Reaching, 2. Uneven terrain navigation using the Reinforce algorithms. The Unity simulator includes diverse terrains and elevations for training and testing. We restrict ourselves to a $100m\times 100m$ region during training. We used identical policy distribution parameters and neural network architectures with all four algorithms during training and evaluation. The policies are trained using Pytorch in the simulator equipped with a Clearpath Husky robot model, and a ROS Melodic platform. The training simulations are executed on a workstation with an Intel Xeon 3.6 GHz CPU and an Nvidia Titan GPU.
@@
Our agent is a differential drive robot with a two-dimensional continuous action space $a=(v,\omega)$ (i.e. linear and angular velocities). The simulator includes even and uneven outdoor terrains and the state space varies for the two navigation tasks and is obtained in real-time from the simulated robot’s odometry and IMU sensors. The state inputs used in the two scenarios are,
@@
  • $d_{goal}\in[0,1]$- Normalized distance (w.r.t the initial straight-line distance) between the robot and its goal;
  • $\alpha_{goal}\in[0,\pi]$- Angle between the robot’s heading direction and the goal;
  • $(\phi,\theta)\in[0,\pi]$- Roll and pitch of robot, respectively.
@@
where, $\mathcal{N}$ denotes Gaussian distribution with variance $\sigma$ and $\beta$ is the amplitude parameter. We set the variance $\sigma=0.2$ to ensure that the $r_{dist}$ only includes two narrow reward peaks near the goal and half-way from the goal. Similarly, $r_{elev}$ is defined as $r_{elev}=\mathbb{I}_{\{|\phi|\geq\pi/6\}}\cup\mathbb{I}_{\{|\theta|\geq\pi/6\}}$. Our reward definitions are significantly sparse compared to the traditional dense rewards used in the literature [41].
(a) Even terrain navigation: The robot is placed in an obstacle-free even terrain environment and random goals in the range of 10-15 meters away from the starting position given at each episode. The state space is $s=[d_{goal},\alpha_{goal}]$ and the overall reward formulation $r_{even}=r_{dist}+r_{head}$.
(b) Uneven terrain navigation: The robot is placed in an obstacle-free, highly uneven terrain environment. To account for this complexity, we input the roll and pitch of the angle of the robot into the state space, $s=[d_{goal},\alpha_{goal},\phi,\theta]$. The objective is to navigate the robot to a goal location along relatively even terrain regions to avoid possible robot flip-overs. The reward is $r_{uneven}=r_{dist}+r_{head}-r_{elev}$.
@@
CCE (Ours) 72 16.28 1.985
Table I: Navigation performance of the policies trained for even and uneven terrain navigation tasks with limited samples. SR denotes success rate, PL denotes average path length, and EC denotes elevation cost. Each method is tested for 100 trials in the outdoor simulator and the average values are reported in the table. Test environments are different from training environments.